OpenStudy (anonymous):

Determine the absolute value of -12|37-69| ?

jimthompson5910 (jim_thompson5910):

-12|37-69| -12|-32| -12(32) -384 So -12|37-69| = -384

OpenStudy (anonymous):

-12|37-69| =-12 ==> -12|-32| =-12(32) =-348